The hypotenuse of a right triangle is 1 in longer than the longer leg. The shorter leg measures 7 in less than the longer leg. Find the measure of the longer leg of the triangle.
Answer by mananth(16946)

longer leg =x
shorter leg = x-7
hypotenuse = x+1
(x+1)^2= x^2+(x-7)^2
x^2+2x+1=x^2+x^2-14x+49
x^2-16x+48=0
x^2-12x-4x+48=0
x(x-12)-4(x-12)=0
(x-12)(x-4)=0
x=12 Or 4
longer leg = 12